The purchase cost of an excavating machine is $200000. After a useful life of 6 years, it is assumed the equipment will be soguys please write neat and clean with proper units thanks
Answer #1

ANSWER

Purchase Cost = $ 200000|- Straight line method: Defericiation (0) = c-s. n 200000 - 50000 216) 0 = 25000 constant of method3) Sencing fund method I= S: (148)2-1 1 150000 X 0.2195 (1+0.217561 J 32625 2.25 1 = 14500 $ 1 yearArl year Assuming the equipment used 1600 kr per hour cost 14500 1600 = 9.0625 $lar 9.0625 $ her so per cost -

================

DEAR STUDENT,

IF YOU HAVE ANY QUERY PLEASE ASK ME IN THE COMMENT BOX,I AM HERE TO HELP YOU.PLEASE GIVE ME POSITIVE RATING..

****************THANK YOU****************